OpenStudy (volleyballlover55):

Consider the following scenario describing the residents of Charlestown: The list pairs resident's ages with their zip codes. Does this scenario represent a function? No, because people who are the same age can have different zip codes No, because people who have the same zip code can be different ages Yes, because people who are the same age can have different zip codes Yes, because people who have the same zip code can be different ages

OpenStudy (volleyballlover55):

@calculusxy

OpenStudy (calculusxy):

Zip Codes represent the y-values and the ages represent the x-values. Function = a domain needs to have specific ranges If you think about it, it would be a function if people who would have the different ages, x, would have same zip code, y. When graphed it would be like:  This would pass the vertical line test

OpenStudy (volleyballlover55):

so C?

OpenStudy (calculusxy):

I think maybe B. Although I am 100% sure.

OpenStudy (volleyballlover55):

its b for sure?

OpenStudy (calculusxy):

As I said, I am not 100% sure.

OpenStudy (volleyballlover55):

oh, ok.. next! Given the relation y = –4x2 – 8, if the input is 2, what is the output? 8 –8 –16 –24

OpenStudy (calculusxy):

But you can try it out if you'd like

OpenStudy (calculusxy):

I am assuming that the function is \(y = -4x^2 - 8\) Substitute 2 for \(x\) \( y = -4(2)^2 - 8\) What do you get?

OpenStudy (calculusxy):

@volleyballlover55 I have to leave in like a minute :\

OpenStudy (volleyballlover55):

is it -8?

OpenStudy (smartnerd1111):

If you need to put an nenrt do ^c

OpenStudy (calculusxy):

No. \( y = -4(2)^2 - 8\) \( y = -4(4) - 8\) \( y = -16 - 8\) \( y = -24\)

OpenStudy (volleyballlover55):

oh D!

OpenStudy (volleyballlover55):

hey @smartnerd1111 since @calculusxy , can you help me with some more please? :)

OpenStudy (smartnerd1111):

Well i am ow rightn maygbe in 10 min.n

OpenStudy (volleyballlover55):

your leaving in 10 min?

OpenStudy (volleyballlover55):

i just have 3 more questions :)

OpenStudy (smartnerd1111):

I am coming in ten min

OpenStudy (volleyballlover55):

oh ok! but please come back :)

OpenStudy (volleyballlover55):

come to my new post! :)